Use the appropriate normal distribution to approximate the resulting binomial distribution A coin is weighted so that the probability of obtaining a head in a single tos:s is 0.6. If the coin is tossed 30 times, find the following probabilities. (Round your answers to four decimal places.) (a) fewer than 15 heads (b) between 15 and 19 heads, inclusive (c) more than 22 heads
Answer #1

Let x =number of heads

given n = 30 , p = 0.6 ,then q = 1 - p = 0.4

consider normal approximation .

then mu = np = 30 x 0.6 = 18

and sigma = Vnpq = 30 x 0.6x 0.4 = 2.68

ie x ~ N (mu , sigma ) with mu = 18 and   sigma = 2.68

Then ,lke coer hap 1s heads S18 112 0 oto -112 Prea 0 . I314 二 5 The prob ahilly that beloe s oad It haads 19 IS 11-18 2.6 8 -00-1.12 o o.31 1-12 o 37 = O. 8686-0.3551 0.5129

C) The probabil: ty mme 022 Loads ( χ > 22 22 -18 2 68 o 1.49 = 0.0681